The Twin Tower credit card company charges you an effective annual rate of 15.90 percent on your credit card. What is the quoted annual percentage rate on your account if the interest is compounded monthly?

A) 14.11%

B) 13.14%

C) 15.90%

D) 18.59%

E) 14.85%
Answer #1

EAR=(1+APR/m)^m-1
where m=compounding periods

0.159=(1+APR/12)^12-1

(1+0.159)=(1+APR/12)^12

(1.159)^(1/12)=1+APR/12

APR=[(1.159)^(1/12)-1]*12

=14.85%(Approx).